# Deep learning for predicting the occurrence of tipping points

### v1

@@ Discussion
Predicting the occurrence of tipping points based on time series is a challenging problem. In this work, we develop a deep learning algorithm that exploits information of normal form from the training systems for tipping points prediction. This algorithm can deal with both regularly-sampled and irregularly-sampled time series. Our results show that the DL algorithm not only outperforms traditional methods in the accuracy of tipping points prediction for regularly-sampled time series, but also achieve accurate prediction for irregularly-sampled time series. Our results pave the way to make effective strategies to prevent and prepare tipping points from various real-world systems(47).
@@
There are three major advantages of our work. First, traditional methods for tipping points prediction can only deal with regularly-sampled time series(9, 10, 11, 12). However, the DL algorithm used in this paper can deal with both regularly-sampled time series and irregularly-sampled time series. This is because the bifurcation parameter series input into the 2D CNN layer contains the information about the sequence of distance between adjacent points, which is used for tipping points prediction. Second, the effect of rate-delayed tipping has been ignored in previous methods for tipping points prediction(9, 10, 11, 12). This is due to the limitations of the theory of fast-slow systems used in previous methods. However, we consider the effect of rate-delayed tipping in our work and label the data by using the fact that the quasi-static attractor loses stability at the occurrence of tipping points in our training set (see Methods for details). The DL algorithm trained with such labels is robust against different changing rates of bifurcation parameter (Supplementary Fig. 16). Third, previous methods for tipping points prediction require test data that includes samples taken near the tipping points where pronounced early warning signals are present, particularly for the bifurcation theory based methods(9, 12). However, most of the time series from our experiments were sampled beyond a certain distance from the tipping points, leading to the unsuccessful performance of these algorithms in predicting tipping points shown in this paper. The DL algorithm, in contrast, can accurately identify the location where the tipping points occur with various distance between the end of the time series and the tipping points.
Most previous methods require full information of all interacting variables in the dynamical systems for predicting tipping points, such as the time series of interacting variables(9, 48). However, this information is often not available especially for large systems. Therefore, it is often not feasible to use those methods for predicting tipping points of real-world systems. In this work, we use time series of only a single variable in the dynamical system as prior information for predicting tipping points. Such time series is commonly available from real-world systems, because it is very likely that only one or few variables in the systems are measured(12). The theoretical foundation of our algorithm lies in Takens embedding theorem(13). Specifically, the time series of a single variable preserves dynamical information of the entire system, which can provide sufficient information for tipping points prediction.
@@
Our work raises several problems worthy of future pursuit. First, we have studied the situation of local codimension-one bifurcations in continuous-time dynamical systems in this paper. For future work, we can focus on tipping points prediction in discrete-time dynamical systems, such as period-doubling bifurcation(7) which arise naturally in physiology(49, 50) and ecology(51). It would be interesting to develop a method that can deal with both regularly-sampled and irregularly-sampled time series for systems with period-doubling bifurcation. Second, it would be interesting to investigate tipping points prediction in systems with other types of bifurcation, such as codimension-two bifurcation and global bifurcation(7). Yet, tipping points prediction in systems with codimension-two bifurcation and global bifurcation are more challenging than that in systems with codimension-one bifurcation. Third, one can develop interpretable machine learning models for tipping points prediction by combing dynamical system theory with neural networks(52, 53, 54, 55, 56), which offer an avenue for making safe and reliable high-stakes decisions for policy makers(57).
@@ Methods
$\frac{dx}{dt}=\sum_{i=1}^{10}a_{i}p_{i}(x,y)$ (1)
$\frac{dy}{dt}=\sum_{i=1}^{10}b_{i}p_{i}(x,y),$
@@
$\lambda=max(Re(eigvals(\begin{bmatrix}\frac{\partial(\sum_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial y}\\ \frac{\partial(\sum_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial y}\end{bmatrix}_{(x,y)=(x^{*},y^{*})}))).$ (2)
@@
For each model where a bifurcation is identified, we perform five independent simulations of the model with varying bifurcation parameter from its initial value to a terminal value beyond the bifurcation point given by AUTO-07P. In these five simulations, we set the changing rate of the bifurcation parameter to remain constant within each simulation where the changing rate of the bifurcation parameter is the change in the bifurcation parameter per unit time with a step size of 0.01, but vary across different simulations. The ratio of each changing rate of the bifurcation parameter to the smallest changing rate of the bifurcation parameter in these five simulations is drawn from 1, 2, …, 10. The nonzero changing rate of the bifurcation parameter can lead to the delay in the occurrence of tipping points, which is called rate-delayed tipping(20, 21). Due to rate-delayed tipping, the theoretical bifurcation point given by AUTO-07P is not the accurate location of the tipping point. Therefore, we identify the location of the tipping point by the recovery rate (Eq. 2) and use this identified tipping point as the training label for our supervised deep learning training process (Supplementary Information section 2.3). If no tipping point is detected in a simulation, this simulated data will be discarded. Otherwise, we randomly and irregularly sample $l_{s}$ points from quasi-static attractor time series with noise which is simulated by the variable $x$ of the model and corresponding bifurcation parameter time series between the initial value and the tipping point, where $l_{s}$ is drawn from $Uniform(505,1000)$. Then we select the first 500 points as training data. The obtained training data can make DL algorithm to learn the features of recovery rate from time series data with different changing rates of parameter and varying distances between the end of the time series and the tipping points.
Here we train deep learning algorithm on datasets with white and red noise. The white noise in simulation has the amplitude drawn from a triangular distribution centered at 0.01 with upper and lower bounds 0.0125 and 0.0075 respectively while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient is between -1 and 1. It is worth noting that the simulations shift to new regime before the bifurcation point due to the noise(62, 63). Since the location of this premature transition is stochastical, we still use the location where the quasi-static attractor without noise loses stability as the training label for DL algorithm even it is not accurate. We mathematically illustrate how slight perturbation causes bifurcation to occur earlier and its stochasticity in Supplementary Information section 6.
@@
In this paper we use the 2D CNN-LSTM DL algorithm(64, 65). The 2D CNN-LSTM architecture is shown in Fig. 6. We use a train/validation/test split of 0.95/0.04/0.01 for the DL algorithm, and MSE as the loss function to minimize the difference between the real tipping points (training labels) and the predictions. The DL algorithm is trained on 300,000 instances for 200 epoches. We train ten neural networks and report the performance of the DL algorithm averaged over them. The hyperparameters of the DL model are tuned through random search, and we present the optimal hyperparameters of the DL model in Supplementary Table 1.
Before training our DL algorithm, there are several preprocessing matters(66). The quasi-static attractor time series with noise are detrended using Lowess smoothing with a span of 0.2 to obtain the residual time series used for training. Besides, in order to obtain the robustness of DL algorithm on time series of various lengths, we zero out the left side of the residual time series and corresponding bifurcation parameter time series, with the length drawn from $Uniform(0,250)$. Due to the zeroing process, the DL algorithm can deal with time series even if it is a little short. After the zeroing process, each residual time series is normalized by dividing each time series data point by the average absolute value of the residuals across the non-zero part of the time series. In addition, each bifurcation parameter time series is also normalized: each data point in the time series minus the initial value of the bifurcation parameter time series, and then divided by the distance between the initial and final values of the bifurcation parameter time series, thereby mapping the bifurcation parameter time series to the interval of $[0,1]$. The corresponding label, i.e., the tipping point, is normalized following the same procedure as the normalization of the data point in bifurcation parameter time series. Thus the relative distance of tipping point to the time series in the training set for the DL algorithm (calculated as the tipping points minus the final value of the bifurcation parameter time series, divided by the distance between the initial and final values of the bifurcation parameter time series) is between 0.01 and 2. In other words, the labels in the training set, after normalization, will fall within the interval of $[1.01,3]$.
@@
The simulation of theoretical models with noise is based on the Euler Maruyama method(60) with a step size of 0.01 ($\delta t=0.01$ units of time) unless otherwise stated. The amplitude of white noise $\sigma$ is 0.01 while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient $\phi$ is drawn from $Uniform(-1,1)$. We fix the changing rate of the bifurcation parameter in each simulation of the same theoretical model to ensure that these simulations have the same bifurcation point. The bifurcation point of each simulation is the location where the recovery rate changes from negative value to positive value. Then we sample from these simulations to obtain regularly-sampled model time series and irregularly-sampled model time series for testing. The lengths of these model time series range from 250 to 500, and the relative distances of tipping points to these time series range from 0.01 to 2.
@@
where $x$ is biomass of some population, $k$ is its carrying capacity, $h$ is the harvesting rate, $s$ characterizes the nonlinear dependence of harvesting output on current biomass, $r$ is the intrinsic per capita growth rate of the population, and $\xi(t)$ is a Gaussian white noise process. We use parameter values $r=1$, $k=1$, $s=0.1$ and $h$ increases at a rate of $6.25\times 10^{-4}$. We generate model time series by this equation from eleven initial values of $h$, which are $0,0.02,0.04,...,0.18,0.2$. In this configuration, a fold bifurcation occurs at $h=0.268$ (Supplementary Fig. 9).
@@
where $r$, $c$, and $p$ are the population densities of the resource, consumer, and predator species, respectively. $k$ is a parameter characterizing the environmental capacity of the resource species. $x_{c}$, $y_{c}$, $x_{p}$, $y_{p}$, $r_{0}$, and $c_{0}$ are other parameters in the system. $\xi_{r}(t)$, $\xi_{c}(t)$ and $\xi_{p}(t)$ are independent Gaussian white noise processes. For our simulations, we use parameter values $x_{c}=0.4$, $y_{c}=2.009$, $x_{p}=0.08$, $y_{p}=2.876$, $r_{0}=0.16129$, $c_{0}=0.5$ and $k$ increases at a rate of $1.54\times 10^{-3}$. We generate model time series by these equations from eleven initial values of $k$, which are $0.20,0.22,0.24,...,0.38,0.40$. In this configuration, a Hopf bifurcation occurs at $k=0.480$ (Supplementary Fig. 10).
@@
where $x$ and $y$ are the population densities of the resource, consumer, respectively. $g$ is the intrinsic per capita growth rate of $x$, $k$ is its carrying capacity, $a$ is the attack rate of $y$, $e$ is the conversion factor, $h$ is the handling time, $m$ is the per capita consumer mortality rate, $\xi_{x}(t)$ and $\xi_{y}(t)$ are independent Gaussian white noise processes. We fix the parameter values $r=4$, $k=1.7$, $e=0.5$, $h=0.15$, $m=2$ and $a$ increases at a rate of $8.112\times 10^{-3}$. We generate model time series by these equations from eleven initial values of $a$, which are $0,0.5,1.0,...,4.5,5.0$. In this configuration, the system has a transcritical bifurcation at $a=5.882$ (Supplementary Fig. 11).
@@
To test the fold bifurcationn with red noise, we use a climate model describing temperature of an ocean on a spherical planet subjected to radiative heating(32, 31) which can simulate a transition from a greenhouse to an icehouse Earth(42). The model is given by
@@
where $T$ is the average temperature of ocean, $u$ is relative intensity of solar radiation, $e$ is effective emissivity, $I_{0}$ is solar irradiance, $c$ is a constant thermal inertia, and $a_{p}$ is the planetary albedo. Parameters $a$ and $b$ define a linear feedback between ice and albedo variability and temperature. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. We use $e=0.69$, $I_{0}=71944000$, $c=10^{8}$, $a=2.8$, $b=0.009$, $\rho=0.03$, $\delta t=1$ time unit and $u$ decreases at a rate of $6\times 10^{-7}$. We generate model time series by this equation from eleven initial values of $u$, which are $1.4,1.38,1.36,...,1.22,1.2$. In this configuration, a fold bifurcation occurs at $u=0.962$ (Supplementary Fig. 12).
@@
where $x$ represents the global ice volume, $y$ represents the atmospheric greenhouse gas concentration and $z$ represents a deep ocean temperature. All variables are rescaled to dimensionless form, $p=1$, $q=1.2$, $s=0.8$ are parameters, and $u$ increases at a rate of $8.4\times e^{-6}$. $\eta_{x}(t)$, $\eta_{y}(t)$ and $\eta_{z}(t)$ are independent red noise processes with lag-1 autoregressive coefficient $\phi$. We generate model time series by these equations from eleven initial values of $u$, which are $0,0.03,0.06,...,0.27,0.3$. In this configuration, a Hopf bifurcation occurs at $u=0.355$ (Supplementary Fig. 13).
To test the transcritical bifurcation with red noise, we use the simplified version of TRIFFID dynamic global vegetation model(34, 67). It can be used to simulate the dieback of the Amazon rainforest(67), which is mainly driven by climate change(68). The model is given by
@@
where $V$ is equal to the broadleaf fraction, $G$ is a disturbance coefficient (0.004/year) and $V^{*}$ is either the value of $V$ or 0.1 if $V$ falls below 0.1. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. $P$ is the productivity, in dimensionless area fraction units, it decreases at a rate of $2.4\times 10^{-3}$. We generate model time series by this equation from eleven initial values of $P$, which are $0.90,0.82,0.74,...,0.18,0.10$. A transcritical bifurcation occurs at $P=-0.005$ in this configuration (Supplementary Fig. 14).
@@
In this work, we use three empirical examples in the fields of microbiology, thermoacoustics, and paleoclimatology to evaluate the performance of the DL algorithm.
[…]
@@
The third example is the concentration of metallic elements in sediment cores collected using a piston corer at 33°18.1’N, 33°23.7’E, 1760 m water depth in the eastern Mediterranean Sea(38) where rapid transitions between oxic and anoxic states occurred regularly. Molybdenum (Mo) and uranium (U) are commonly used as key indicators for anoxic events, which signify past climatic or environmental changes that occurred during certain periods(70, 71, 72). Therefore, we select the time series of the concentration of Mo and U in sediment cores for our study. In the analysis of these empirical datasets, we assume that the bifurcation parameter of the system varies uniformly with time, such that age can be served as a substitute for this parameter. The time series of the concentration of Mo in sediment cores consists of 1,147 data points ranging from 160 ka BP to 125 ka BP while the time series of the concentration of U in sediment cores consists of 1,142 data points ranging from 300 ka BP to 240 ka BP.
@@
For detecting early warning signals (EWS) in quasi-static attractor time series with white noise, H. Held and T. Kleinen(9) develop a technique called degenerate fingerprinting using PCA to approximate the critical mode, then estimating the lag-1 autoregressive coefficient of critical mode as an indicator (Supplementary Information section 4.1). Boettner and Boers(73, 10) propose an unbiased estimate of the lag-1 autoregressive coefficient for time series with red noise which we refer to as the BB method (Supplementary Information section 4.2). Florian Grziwotz et al.(12) introduce a EWS robust to limited level of noise, named dynamical eigenvalue (DEV), which is rooted in bifurcation theory of dynamical systems to estimate the dominant eigenvalue of the system (Supplementary Information section 4.3). After obtaining EWS through these three approaches, linear regression or nonlinear fit between EWS and bifurcation parameters is used to make an extrapolation to anticipate the occurrence of tipping points. Based on experimentation, we choose to fit a quadratic curve for extrapolation, which is proved to be more robust than linear regression.
It is worth noting that the required information of detecting early warning signals with degenerate fingerprinting is different from that with our DL algorithm. The degenerate fingerprinting requires information from all variables of the study system to approximate the critical mode, whereas our DL algorithm only requires data from one variable of the study system. Similarly, the DEV also requires data from one variable of the study system. Moreover, the BB method is designed for estimating the lag-1 autoregressive coefficient of one-dimensional system, whereas our DL algorithm is suitable for analyzing high-dimensional system. The DEV can also deal with high-dimensional system.
Null Model
To create a new training set for null model, we shuffle the residual time series in the original training set, completely randomizing the order of data points in the residual time series, while keeping the bifurcation parameter time series unchanged. Then we use the 2D CNN-LSTM DL algorithm to train null model on this dataset. Following random search, the optimal hyperparameters of the null model are presented in Supplementary Table 1. We compare our DL model with this null model on model time series and empirical time series to validate that the DL algorithm has indeed extracted useful information from the data.
@@
All the simulated datasets and the empirical datasets used to test the deep learning algorithm have been deposited in Github (https://github.com/zhugchzo/dl_occurrence_tipping) and the training set used to train the deep learning algorithm have been deposited in Zenodo (https://zenodo.org/records/12725553).

### v2

@@ Discussion
Predicting the occurrence of tipping points based on time series is a challenging problem. In this work, we develop a deep learning algorithm that exploits information of normal form from the training systems for tipping points prediction. This algorithm can deal with both regularly-sampled and irregularly-sampled time series by using the time series sampled from a single variable of a system. Our results show that the DL algorithm not only outperforms traditional methods in the accuracy of tipping points prediction for regularly-sampled time series, but also achieve accurate prediction for irregularly-sampled time series. Our results pave the way to make effective strategies to prevent and prepare tipping points from various real-world systems(47).
@@
There are three major advantages of our work. First, traditional methods for tipping points prediction can only deal with regularly-sampled time series(9, 10, 11, 12). However, the DL algorithm used in this paper can deal with both regularly-sampled time series and irregularly-sampled time series. This is because the convolution kernels in the 2D CNN layer can extract features of the reconstructed system from segments of irregularly-sampled time series (Supplementary information Note S3), which is used for tipping points prediction in the LSTM layer. Second, most existing methods(9, 48, 25, 26, 27) require information of multiple interacting variables in the study system for predicting tipping points. However, based on the embedding theorem for irregular sampling(28), the DL algorithm only requires the time series sampled from a single variable of the study system. Third, the effect of rate-delayed tipping has been ignored in previous methods for tipping points prediction(9, 10, 11, 12). This is due to the limitations of the theory of fast-slow systems used in previous methods. However, we consider the effect of rate-delayed tipping in our work and label the data by using the fact that the quasi-static attractor loses stability at the occurrence of tipping points in our training set (see Methods for details). The DL algorithm trained with such labels is robust against different changing rates of bifurcation parameter (Supplementary information Figs. S5-S7).
The embedding theorem for irregular sampling(28) requires that the dimension $d$ of the system reconstructed by delay embedding must be larger than twice the dimension $m$ of the study system (Supplementary information Note S3). Therefore, the length of convolution kernels used in the CNN layer of our DL algorithm is required to be larger than $2m$. However, due to the time-varying nonstationarity of dynamical system approaching a bifurcation, the features extracted from shorter convolutional kernel should contain much more dynamical information of the system, compared to those extracted from longer convolutional kernel(49). Therefore, our DL algorithm may not perform well in predicting tipping points in high-dimensional systems. Another limitation of our DL algorithm is that it requires the existence of tipping points as prior knowledge. If there is no tipping point in the system, the output of the DL algorithm is meaningless. Therefore, several methods such as DEV(12) or some DL classifiers(18, 22, 23, 24) should be used to determine whether a system is approaching a tipping point first. Then our DL algorithm can be applied to predict the specific location of a tipping point if it exists.
Fig. 6. The performance of DL algorithm in predicting tipping points on irregularly-sampled empirical time series. The dots denote the DL predictions, and the short vertical lines denote the initial points of the time series data used for prediction. We connect them with lines. (A1-A3) Cyanobacterial population undergoes a fold bifurcation, which is measured by light attenuation coefficient. The fold bifurcation occurs at 1091 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$, while the predictions by our DL algorithm are between 1074 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ and 1106 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. (B1-B3) The thermoacoustic system undergoes a Hopf bifurcation under the 40 mV/s changing rate of the control parameter voltage. The Hopf bifurcation occurs at 1.76 V, while the predictions by our DL algorithm are between 1.71 V and 1.79 V. We compared the performance of the DL algorithm (red) with degenerate fingerprinting (blue) and DEV (green). We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting and DEV.
To analyze whether the DL algorithm uses the recovery rate in the normal form for prediction, we designed two control experiments. First, we trained three DL models on three datasets, each consisting solely of time series with fold, Hopf, and transcritical bifurcation, respectively. Then we applied these three DL models on irregularly-sampled model time series from six theoretical models. We found that the DL model trained on dataset containing the features of a bifurcation perform better in predicting tipping points of test time series with that bifurcation, compared to another two trained on datasets without the corresponding normal form features (Supplementary information Fig. S8). These results suggest that the DL algorithm can extract the features of normal form. Second, we trained two DL models on two datasets, each consisting solely of time series with supercritical and subcritical pitchfork bifurcation, respectively (Supplementary information Note S7). Then we applied these two DL models on irregularly-sampled model time series from a model with supercritical pitchfork, where both models achieved similar performance in predicting the tipping points (Supplementary information Fig. S9). The normal forms of supercritical and subcritical pitchfork bifurcations exhibit the same relation between the recovery rate and the bifurcation parameter, and only differ in the cubic term. Thus, the DL model trained on time series with subcritical pitchfork bifurcation can be used to predict tipping points of time series with supercritical pitchfork bifurcation, which suggest that the DL algorithm can extract the features of the recovery rate in normal form.
@@
Our work raises several problems worthy of future pursuit. First, we have studied the situation of local codimension-one bifurcations in continuous-time dynamical systems in this paper. For future work, we can focus on tipping points prediction in discrete-time dynamical systems, such as period-doubling bifurcation(7) which arise naturally in physiology(50, 51) and ecology(52). It would be interesting to develop a method that can deal with both regularly-sampled and irregularly-sampled time series for systems with period-doubling bifurcation. Second, it would be interesting to investigate tipping points prediction in systems with other types of bifurcation, such as codimension-two bifurcation and global bifurcation(7). Yet, tipping points prediction in systems with codimension-two bifurcation and global bifurcation are more challenging than that in systems with codimension-one bifurcation. Third, one can develop interpretable machine learning models for tipping points prediction by combing dynamical system theory with neural networks(53, 54, 55, 56, 57), which offer an avenue for making safe and reliable high-stakes decisions for policy makers(58).
@@ Methods
$\frac{dx}{dt}=\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y)$ (1)
$\frac{dy}{dt}=\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y),$
@@
$\lambda=max(Re(eigvals(\begin{bmatrix}\frac{\partial(\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial y}\\ \frac{\partial(\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial y}\end{bmatrix}_{(x,y)=(x^{*},y^{*})}))).$ (2)
@@
For each model where a bifurcation is identified, we perform five independent simulations of the model with varying bifurcation parameter from its initial value to a terminal value beyond the bifurcation point given by AUTO-07P. In these five simulations, we set the changing rate of the bifurcation parameter to remain constant within each simulation where the changing rate of the bifurcation parameter is the change in the bifurcation parameter per unit time with a step size of 0.01, but vary across different simulations. The ratio of each changing rate of the bifurcation parameter to the smallest changing rate of the bifurcation parameter in these five simulations is drawn from $1,2,\dots,10$. The nonzero changing rate of the bifurcation parameter can lead to the delay in the occurrence of tipping points, which is called rate-delayed tipping(20, 21). Due to rate-delayed tipping, the theoretical bifurcation point given by AUTO-07P is not the accurate location of the tipping point. Therefore, we identify the location of the tipping point by the recovery rate (Eq. 2) and use this identified tipping point as the training label for our supervised deep learning training process (Supplementary information Note S2). If no tipping point is detected in a simulation, this simulated data will be discarded. Otherwise, we randomly and irregularly sample $l_{s}$ points from quasi-static attractor time series with noise which is simulated by the variable $x$ of the model and corresponding bifurcation parameter time series between the initial value and the tipping point, where $l_{s}$ is drawn from $Uniform(505,1000)$. Then we select the first 500 points as training data. The obtained training data can make DL algorithm to learn the features of recovery rate from time series data with different changing rates of parameter and varying distances between the end of the time series and the tipping points.
Here we train deep learning algorithm on datasets with white and red noise. The white noise in simulation has the amplitude drawn from a triangular distribution centered at 0.01 with upper and lower bounds 0.0125 and 0.0075 respectively while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient is between -1 and 1. It is worth noting that the simulations shift to new regime before the bifurcation point due to the noise(63, 64). Since the location of this premature transition is stochastical, we still use the location where the quasi-static attractor without noise loses stability as the training label for DL algorithm even it is not accurate. We mathematically illustrate how slight perturbation causes bifurcation to occur earlier and its stochasticity in Supplementary information Note S6.
@@
In this paper we use the 2D CNN-LSTM DL algorithm(65, 66). The 2D CNN-LSTM architecture is shown in Fig. 1. We use a train/validation/test split of 0.95/0.04/0.01 for the DL algorithm, and MSE as the loss function to minimize the difference between the real tipping points (training labels) and the predictions. The DL algorithm is trained on 300,000 instances for 200 epoches. We train ten neural networks and report the performance of the DL algorithm averaged over them. The hyperparameters of the DL model are tuned through random search, and we present the optimal hyperparameters of the DL model in Supplementary information Table S1.
Before training our DL algorithm, there are several preprocessing matters(67). The quasi-static attractor time series with noise are detrended using Lowess smoothing with a span of 0.2 to obtain the residual time series used for training(68). Besides, in order to obtain the robustness of DL algorithm on time series of various lengths, we zero out the left side of the residual time series and corresponding bifurcation parameter time series, with the length drawn from $Uniform(0,250)$. Due to the zeroing process, the DL algorithm can deal with time series even if it is a little short. After the zeroing process, each residual time series is normalized by dividing each time series data point by the average absolute value of the residuals across the non-zero part of the time series. In addition, each bifurcation parameter time series is also normalized: each data point in the time series minus the initial value of the bifurcation parameter time series, and then divided by the distance between the initial and final values of the bifurcation parameter time series, thereby mapping the bifurcation parameter time series to the interval of $[0,1]$. The corresponding label, i.e., the tipping point, is normalized following the same procedure as the normalization of the data point in bifurcation parameter time series. Thus the relative distance of tipping point to the time series in the training set for the DL algorithm (calculated as the tipping points minus the final value of the bifurcation parameter time series, divided by the distance between the initial and final values of the bifurcation parameter time series) is between 0.01 and 2. In other words, the labels in the training set, after normalization, will fall within the interval of $[1.01,3]$.
@@
The simulation of theoretical models with noise is based on the Euler Maruyama method(61) with a step size of 0.01 ($\delta t=0.01$ units of time) unless otherwise stated. The amplitude of white noise $\sigma$ is 0.01 while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient $\phi$ is drawn from $Uniform(-1,1)$. We set five different changing rates of the bifurcation parameter in simulations of each theoretical model. The bifurcation point of each simulation is the location where the recovery rate changes from negative value to positive value. Then we sample from these simulations to obtain regularly-sampled model time series and irregularly-sampled model time series for testing. The lengths of these model time series range from 250 to 500, and the relative distances of tipping points to these time series range from 0.01 to 2.
@@
where $x$ represents biomass of some population, $k$ is its carrying capacity, $h$ is the harvesting rate, $s$ characterizes the nonlinear relationship between harvesting output and current biomass, $r$ is the intrinsic per capita growth rate of the population, and $\xi(t)$ is a Gaussian white noise process. We use parameter values $r=1$, $k=1$, $s=0.1$ and $h$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by this equation from eleven initial values of $h$, which are $0,0.02,0.04,\dots,0.18,0.2$. In this configuration, the fold bifurcation occurs in $h\in[0.2713,0.2926]$.
@@
where $r$, $c$, and $p$ represent the population densities of the resource, consumer, and predator species, respectively. $k$ characterizes the environmental capacity of the resource species. $x_{c}$, $y_{c}$, $x_{p}$, $y_{p}$, $r_{0}$, and $c_{0}$ are other parameters in the system. $\xi_{r}(t)$, $\xi_{c}(t)$ and $\xi_{p}(t)$ are independent Gaussian white noise processes. For our simulations, we use parameter values $x_{c}=0.4$, $y_{c}=2.009$, $x_{p}=0.08$, $y_{p}=2.876$, $r_{0}=0.16129$, $c_{0}=0.5$ and $k$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by these equations from eleven initial values of $k$, which are $0.20,0.22,0.24,\dots,0.38,0.40$. In this configuration, the Hopf bifurcation occurs in $k\in[0.4656,0.4839]$.
@@
where $x$ and $y$ represent the population densities of the resource, consumer, respectively. $g$ is the intrinsic per capita growth rate of $x$, $k$ is its carrying capacity, $a$ is the attack rate of $y$, $e$ is the conversion factor, $h$ is the handling time, $m$ is the per capita consumer mortality rate, $\xi_{x}(t)$ and $\xi_{y}(t)$ are independent Gaussian white noise processes. We fix the parameter values $r=4$, $k=1.7$, $e=0.5$, $h=0.15$, $m=2$ and $a$ increases at rates of $1\times 10^{-4},2\times 10^{-4},\dots,5\times 10^{-4}$. We generate model time series by these equations from eleven initial values of $a$, which are $0,0.5,1.0,\dots,4.5,5.0$. In this configuration, the transcritical bifurcation in $a\in[5.8819,5.8823]$.
@@
To test the fold bifurcation with red noise, we use a climate model describing temperature of an ocean on a spherical planet subjected to radiative heating(33, 32) which can simulate a transition from a greenhouse to an icehouse Earth(42). The model is given by
@@
where $T$ represents the average temperature of ocean, $u$ is relative intensity of solar radiation, $e$ is effective emissivity, $I_{0}$ is solar irradiance, $c$ is a constant thermal inertia, and $a_{p}$ is the planetary albedo. $a$ and $b$ are parameters which define a linear feedback between ice and albedo variability and temperature. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. We use $e=0.69$, $I_{0}=71944000$, $c=10^{8}$, $a=2.8$, $b=0.009$, $\rho=0.03$, $\delta t=1$ time unit and $u$ decreases at rates of $5\times 10^{-7},6\times 10^{-7},\dots,9\times 10^{-7}$. We generate model time series by this equation from eleven initial values of $u$, which are $1.4,1.38,1.36,\dots,1.22,1.2$. In this configuration, the fold bifurcation occurs in $u\in[0.9596,0.9631]$.
@@
where $x$, $y$ and $z$ represent the global ice volume, the atmospheric greenhouse gas concentration and the deep ocean temperature, respectively. All variables are rescaled to dimensionless form, $p=1$, $q=1.2$, $s=0.8$ are parameters, and $u$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. $\eta_{x}(t)$, $\eta_{y}(t)$ and $\eta_{z}(t)$ are independent red noise processes with lag-1 autoregressive coefficient $\phi$. We generate model time series by these equations from eleven initial values of $u$, which are $0,0.03,0.06,\dots,0.27,0.3$. In this configuration, the Hopf bifurcation occurs in $u\in[0.3546,0.3547]$.
To test the transcritical bifurcation with red noise, we use the simplified version of TRIFFID dynamic global vegetation model(35, 69). It can be used to simulate the dieback of the Amazon rainforest(69), which is mainly driven by climate change(70). The model is given by
@@
where $V$ represents the broadleaf fraction, $G$ is a disturbance coefficient (0.004/year) and $V^{*}$ is either the value of $V$ or 0.1 if $V$ falls below 0.1. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. $P$ is the productivity, in dimensionless area fraction units, it decreases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by this equation from eleven initial values of $P$, which are $0.90,0.82,0.74,\dots,0.18,0.10$. In this configuration, the transcritical bifurcation occurs in $P\in[-0.0061,-0.0046]$.
@@
In this work, we use two empirical examples in the fields of microbiology and thermoacoustics to evaluate the performance of the DL algorithm.
In the first example, photo-inhibition drives cyanobacterial population to a fold bifurcation when a critical light level is exceeded(38). In this system, bifurcation parameter light irradiance starts at 477 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ and is increased in steps of 23 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ each day. The time series includes 7,784 data points spanning overall 28.86 days with time interval equal to 0.0035 day (5 min).
@@
For detecting early warning signals (EWS) in quasi-static attractor time series with white noise, H. Held and T. Kleinen(9) develop a technique called degenerate fingerprinting using PCA to approximate the critical mode, then estimating the lag-1 autoregressive coefficient of critical mode as an indicator (Supplementary information Note S5). Boettner and Boers(72, 10) propose an unbiased estimate of the lag-1 autoregressive coefficient for time series with red noise which we refer to as the BB method (Supplementary information Note S5). Florian Grziwotz et al.(12) introduce a EWS robust to limited level of noise, named dynamical eigenvalue (DEV), which is rooted in bifurcation theory of dynamical systems to estimate the dominant eigenvalue of the system (Supplementary information Note S5). After obtaining EWS through these three approaches, linear regression or nonlinear fit between EWS and bifurcation parameters is used to make an extrapolation to anticipate the occurrence of tipping points. In our comparative experiment, we employ linear regression or fit a quadratic curve for extrapolation, and then we select the estimate with the better performance.
It is worth noting that the required information of detecting early warning signals with degenerate fingerprinting is different from that with our DL algorithm. The degenerate fingerprinting requires information from all variables of the study system to approximate the critical mode, whereas our DL algorithm only requires data from one variable of the study system. Similarly, the DEV also requires data from one variable of the study system. Moreover, the BB method is designed for estimating the lag-1 autoregressive coefficient of one-dimensional system, whereas our DL algorithm is suitable for analyzing multidimensional system. The DEV can also deal with multidimensional system.
@@
All the simulated datasets and the empirical datasets used to test the deep learning algorithm have been deposited in Github (https://github.com/zhugchzo/dl_occurrence_tipping) and the training set used to train the deep learning algorithm have been deposited in Zenodo (https://zenodo.org/records/13894933).